OpenStudy (anonymous):

pls help.... prove: tan 70° = tan 20° + 2tan 50°

OpenStudy (anonymous):

I think the ans required is ......... tan(70-20)=tan50 tan70-tan20/1+tan20 tan70=tan50 tan20 tan70=cot70 tan70=1 tan70-tan20/2=tan50 tan70-tan20=2tan50 tan70=tan20+2tan50

OpenStudy (anonymous):

if the ans is satisfactory, pls click GOOD ANSWER button ( :))

OpenStudy (anonymous):

tnx a lot.... :)